OpenStudy (anonymous):

I'm confused with a couple problems I want to answer before a test in an hour. How do you solve 81-1x4 the is squared...

OpenStudy (aum):

question not clear. 1x^4 ? "the is squared" ??

OpenStudy (aum):

Can you post a screenshot of the question?

OpenStudy (anonymous):

\[81-1x^{4}\]

OpenStudy (anonymous):

\[81-1x ^{4}\] Could you help me with it?

OpenStudy (aum):

Solving requires an equation. There is no equation here. We can factor it if that is what is asked. But since you are not posting the full question it is hard to answer the question.

OpenStudy (anonymous):

Sorry it is factoring, im not sure how to break it down.

OpenStudy (aum):

I will assume they are asking you to factor it: \[ 81 - 1x^4 = 81 - x^4 = 9^2 - (x^2)^2 = (9+x^2)(9-x^2) \\ 9 - x^2 = 3^2 - x^2 = (3 + x)(3-x) \\ 81 - 1x^4 = (9+x^2)(9-x^2) = (9+x^2)(3+x)(3-x) \]

OpenStudy (aum):

\(a^2-b^2 = (a+b)(a-b) \)

OpenStudy (anonymous):

One more problem and that was \[1x ^{3}+64\] If you do not mind helping..

OpenStudy (aum):

\[\large a^3 + b^3 = (a+b)(a^2 - ab + b^2) \\ \large 1x^3 + 64 = x^3 + 64 = x^3 + 4^3 = (x+4)(x^2 - 4x + 16) \]

OpenStudy (anonymous):

thank you it is really appreciated!!!

OpenStudy (aum):

You are welcome.